Help! Please show work and explain.. please
Done Lesson 1 Homework Approximately 1.5 million full-time, first-year students enrolled in colleges and universities in 2013. What do they plan to study? Suppose these are the data on the percents of first year students who plan to major in several discipline areas Percent of Students 10.9 10.8 14,7 5.8 12 3.8 15.3 12.2 0.9 13.4 Field of Study Arts and humanities Biological sciences Business Education Engineering Physical sciences Professional Social sciences Technical Other majors and undeclared Roundoff error- %.
Answer #1

The sum of all percent of students should sum to 100.

Sum of all percent of students = 10.9 + 10.8 + 14.7 + 5.8 + 12 + 3.8 + 15.3 + 12.2 + 0.9 + 13.4 = 99.8

Round off error = 100 - 99.8 = 0.2 %